Circular No. 6722

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970815 T. Murakami, Y. Ueda, M. Ishida, and R. Fujimoto, Institute of Space

and Astronautical Science; and A. Yoshida and N. Kawai, Institute of Physical and Chemical Research (RIKEN), report on behalf of the ASCA team: “ASCA observed the error box of GRB 970815 following the detection by ASM/RXTE (IAUC 6718) from Aug. 18.68 to 19.80 UT (net exposures of 40 000 and 46 000 s for the SIS and GIS detectors, respectively). Analysis of the whole data set reveals no source above 1 × 10^{−13} erg cm^{−2} s^{−1} (2– 10 keV) within the ASM/RXTE error box. A nearby x-ray source was seen in both the SIS and GIS fields-of-view, slightly outside the ASM/RXTE error box, at α = 16^{h}06^{m}54^{s}, δ = +81^{o}30′34′′ (1′ error radius); its flux was 3 × 10^{−13} erg cm^{−2} s^{−1} at 2–10 keV, and it seemed stable during the ASCA observation.”

GRB 970616 J. Greiner, Astrophysikalisches Institut Potsdam; J. van Paradijs, Uni-

versity of Amsterdam and University of Alabama at Huntsville; F. M. Marshall, Goddard Space Flight Center; K. Hurley, University of California at Berkeley; C. Robinson, Marshall Space Flight Center; and J. Siebert, Max-Planck-Institut für Extraterrestrische Physik, Garching, report: “A 22 000s ROSAT observation of the RXTE error box of GRB 970616 (IAUC 6683) during June 23–25 reveals eleven faint x-ray sources within the HRI fieldof-view. None of these coincides with the two optical variables reported on IAUC 6687 and 6690, and three are within the IPN/RXTE error box (IAUC 6687): α = 1^{h}19^{m}03^{s}.0, δ = −5^{o}31′35′′; 1^{h}18^{m}50^{s}.3, −5^{o}25′58′′; 1^{h}18^{m}17^{s}.3, −5^{o}12′48′′ (equinox 2000.0). The second of these sources coincides with the first ASCA variable source (IAUC 6687). The unabsorbed ROSAT flux (using a power-law model with photon index 2 and galactic absorbing column N_{H} = 4 × 10^{20} cm^{−2}) is 1.4 × 10^{−14} erg cm^{−2} s^{−1} in the range 0.5–2 keV (i.e., a factor of 5 lower than the flux measured by ASCA on June 20–21, using the same model assumptions and energy range). The 3σ upper limit of the flux at the position of the fourth ASCA source is a factor of 7 lower than that of the ASCA detection. With two fading x-ray sources within the RXTE/IPN error box, an association to the GRB remains open (an image

can be viewed at http://www.aip.de:8080/~jcg/grb970616.html).”

1997 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green